use log(4)3 is about 0.7925 and log(4)4 is about 1 to approximate the value of the expression log(4)96
Answer by stanbon(75887)

use log(4)3 is about 0.7925 and log(4)4 is 1 to approximate the value of the expression log(4)96
---------------------------------
96 = 4*24 = 4*4*6 = 4*4*2*3 = 2^5*3
---
Note: You need a value of log4(2); It is 1/2
----------------------------------------------------
So, log4(96) = log4(2^5*3)
= 5log4(2) + log4(3)
= 5*(1/2) + 0.7925
= 3.2925
===============
Cheers,
Stan H.
===============